On Memory Models and Cache Management for Shared-Memory Multiprocessors

Computation Structures Group Memo 363 March 16, 1995

Jack B. Dennis and Guang R. Gao

This memo is a paper prepared for the IEEE Symposium on Parallel and Distributed Processing, San Antonio, October 1995

The research reported in this docment was performed, in part, using facilities of the Laboratory for Computer Science of the Massachusetts Institute of Technology. Funding for the Laboratory is provided in part by the Advanced Research Projects Agency of the Department of Defense under the Office of Naval Research contract N00014-92-J-1310.